A 15g bullet traveling at 500m/s strikes an 0.8kg block of wood that is balanced on the table edge 0.8m above the ground. If the bullet buries itself in the block . Find the distance D at wich the block hits the floor.

1
Expert's answer
2021-12-10T11:31:04-0500

The given parameters:

mass of the bullet, m1 = 0.015 kg

speed of the bullet, u1 = 500 m/s

mass of block wood, m2 = 0.8 kg

height of the table, h = 0.8 m


The final velocity of the bullet-block system after the collision is calculated by applying the principle of conservation of linear momentum:

"m_1u_1=v(m_1+m_2)"


"v=\frac{m_1u_1}{m_1+m_2}=\frac{0.015\cdot500}{0.015+0.8}=9.2" m/s


The time taken for the bullet-block system to fall to the floor after collision is calculated as follows:

"h=v_{0y}t+gt^2\/2=gt^2\/2"

"t=\sqrt{2h/g}=\sqrt{2\cdot0.8/g}=0.4" s


The horizontal distance where the block hits the floor is calculated as follows:

"x=vt=9.2\cdot0.4=3.68" m